OpenStudy (anonymous):

Find the product. x5 · x4 · x3

OpenStudy (saifoo.khan):

\[\Huge x^{5+4+3}\]

OpenStudy (anonymous):

x^12

OpenStudy (anonymous):

they are all like terms so you can combine them all. multiply them all together. And you end up with 60x

myininaya (myininaya):

\[x^5 \cdot x^4 \cdot x^3=x^{5+4+3}\]

myininaya (myininaya):

\[x(5) \cdot x(4) \cdot x(3)=x \cdot x \cdot x \cdot 5 \cdot 4 \cdot 3=x^3(60)=60x^3\]

OpenStudy (anonymous):

lolz i was so confused for a moment there. in my head i was like " how did you get his answer" and then i realized you were doing it as if they aren't exponents

myininaya (myininaya):

yes lol

myininaya (myininaya):

but still in either case we can't get mamy's answer